### v1

# Latent Ewald summation for machine learning of long-range interactions

Abstract
Machine learning interatomic potentials (MLIPs) often neglect long-range interactions, such as electrostatic and dispersion forces. In this work, we introduce a straightforward and efficient method to account for long-range interactions by learning a latent variable from local atomic descriptors and applying an Ewald summation to this variable. We demonstrate that in systems including charged, polar, or apolar molecular dimers, bulk water, and water-vapor interface, standard short-ranged MLIPs can lead to unphysical predictions even when employing message passing. The long-range models effectively eliminate these artifacts, with only about twice the computational cost of short-range MLIPs.
I Introduction
Machine learning interatomic potentials (MLIPs) can learn from reference quantum mechanical calculations and then predict the energy and forces of atomic configurations quickly, thus allowing for a more accurate and comprehensive exploration of material and molecular properties at scale Keith et al. (2021); Unke et al. (2021a). Most state-of-the-art MLIP methods use a short-range approximation: the effective potential energy surface experienced by one atom is determined by its atomic neighborhood. This approximation implies that the total energy is the sum of atomic contributions, which also makes the MLIPs scale linearly with system size.
The short-range MLIPs, however, neglect all kinds of long-range interactions such as Coulomb and dispersion. Although short-range potentials may be sufficient for describing most properties of homogeneous bulk systems Yue et al. (2021), they may fail for liquid-vapor interfaces Niblett et al. (2021), dielectric response Rodgers and Weeks (2008); Cox (2020), dilute ionic solutions with Debye-Hückel screening, and interactions between gas phase molecules Huguenin-Dumittan et al. (2023).
There has been a continuous effort to incorporate long-range interactions into MLIPs. One can include empirical electrostatics and dispersion baseline corrections Niblett et al. (2021); Lee et al. (2022); Unke et al. (2021b), but for many systems, such baseline is not readily available. Another option is to predict effective partial charges to each atom, which are then used to calculate long-range electrostatics Unke and Meuwly (2019); Ko et al. (2021); Gao and Remsing (2022); Sifain et al. (2018); Gong et al. (2024); Shaidu et al. (2024). For example, the fourth-generation high-dimensional neural network potential (4G-HDNNPs) Ko et al. (2021) predicts the electronegativities of each nucleus, and then use a charge equilibration scheme Rappe and Goddard III (1991) to assign the charges. 4G-HDNNPs are trained directly to reproduce atomic partial charges from reference quantum mechanical calculations, although partial charges are not physical observables and their values depend on the specific partitioning scheme used Sifain et al. (2018). In a similar vein, the deep potential long-range (DPLR) Zhang et al. (2022) learns maximally localized Wannier function centers (MLWFCs) for insulating systems, and the self-consistent field neural network (SCFNN) Gao and Remsing (2022) predicts the electronic response via the position of the MLWFCs. Message passing neural networks (MPNNs) Schütt et al. (2017); Batzner et al. (2022); Deng et al. (2023); Haghighatlari et al. (2022) employ a number of graph convolution layers to communicate information between atoms, thereby capturing long-range interaction up to the local cutoff radius times the number of layers. However, if parts of the system are disconnected on the graph, e.g. two molecules with a distance beyond the cutoff, the message passing scheme does not help. A very interesting approach is the long-distance equivariant (LODE) method Grisafi and Ceriotti (2019); Huguenin-Dumittan et al. (2023), which uses local descriptors to encode the Coulomb and other asymptotic decaying potentials ($1/r^{p}$) around the atoms.
Here, we propose a simple method, the Latent Ewald Summation (LES), for accounting for long-range interactions of atomistic systems. The method is general and can be incorporated into most existing MLIP architectures, including potentials based on local atomic environments (e.g. HDNNP Behler and Parrinello (2007), Gaussian Approximation Potentials (GAP) Bartók et al. (2010), Moment Tensor Potentials (MTPs) Shapeev (2016), atomic cluster expansion (ACE) Drautz (2019)) and MPNN (e.g. NequIP Batzner et al. (2022), MACE Batatia et al. (2022)). In the present work, we combine LES with Cartesian atomic cluster expansion (CACE) MLIP Cheng (2024). After describing the algorithm, we benchmark LES on selected molecular and material systems.
II Theory
For a periodic atomic system, the total potential energy is decomposed into a short-range and a long-range part, i.e. $E=E^{sr}+E^{lr}$. As is standard in most MLIPs, the short-range energy is summed over the atomic contribution of each atom $i$,
$E^{sr}=\sum_{i}E_{\theta}(B_{i}),$ (1)
where $E_{\theta}$ is a multilayer perceptron with parameters $\theta$ that maps the invariant features ($B$) of an atom to its short-range atomic energy. $B$ can be any invariant features used in different MLIP methods, including those based on local atomic environment descriptors such as ACE Drautz (2019), atom centered symmetry functions Behler and Parrinello (2007), smooth overlap of atomic positions (SOAP) Bartók et al. (2013), or any latent invariant features in MPNNs.
For the long-range part, another multilayer perceptron with parameters $\phi$ maps the invariant features of each atom $i$ to a hidden variable, i.e.
$q_{i}=Q_{\phi}(B_{i}).$ (2)
The structure factor $S(\mathbf{k})$ of the hidden variable is defined as
$S(\mathbf{k})=\sum_{i}q_{i}e^{i\mathbf{k}\mathbf{r}_{i}},$ (3)
where $\mathbf{k}=(2\pi n_{x}/L_{x},2\pi n_{y}/L_{y},2\pi n_{z}/L_{z})$ is a reciprocal vector of the orthorhombic cell, and $\mathbf{r}_{i}$ is the Cartesian coordinates of atom $i$. The long-range energy is then obtained using the Ewald summation:
$E^{lr}=\dfrac{1}{V}\sum_{0<k<k_{c}}\dfrac{e^{-\sigma^{2}k^{2}/2}}{k^{2}}|S(\mathbf{k})|^{2},$ (4)
where $\sigma$ is a smearing factor which we typically set to 1 Å, $k=|\mathbf{k}|$ is the magnitude, and $k_{c}$ is the maximum cutoff. $q$ can be multi-dimensional, in which case the total long-range energy is aggregated over contributions from different dimensions of $q$ after the Ewald summation.
The LES method can be interpreted in two ways. First, the hidden variable $q$ is analogous to environment-dependent partial charges on each atom. This implies that the method is at least as expressive as those explicitly based on learning partial atomic charges, as it would yield the same results if $q$ replicated the partial charges. Additionally, $q$ can be multi-dimensional, potentially enhancing expressiveness further. Unlike partial charges, $q$ is not constrained by requirements such as charge neutrality or correct charge magnitudes. Since the Ewald summation of $q$ doesn’t need to correspond to a physical electrostatic potential, Eqn. (4) omits the self-interaction term present in the regular Ewald summation for long-range charge interactions. The second interpretation of LES is as a mechanism that allows atoms far apart in the simulation box to communicate their local information. In this sense, LES is related to the recent Ewald-based long-range message-passing method, which facilitates message exchange between atoms in reciprocal space Kosmala et al. (2023).
III Examples
III.1 Molecular dimers
We benchmark the LES method on the binding curves between dimers of charged (C), polar (P), and apolar (A) relaxed molecules at various separations in a periodic cubic box with a 30 Å edge length. The dataset Huguenin-Dumittan et al. (2023) includes energy and force information calculated using the HSE06 hybrid density functional theory (DFT) with a many-body dispersion correction. We selected one example from each of the six dimer classes (CC, CP, PP, CA, CP, AA), derived from the combination of the three monomer categories. Figure 1 shows a snapshot of each example.
The goal of this benchmark is to evaluate whether the MLIP models can extrapolate dimer interactions at larger separations based on training data from smaller distances. For each molecular pair, the training set consists of 10 configurations with dimer separation distances between approximately 5 Å and 12 Å, and the testing set includes 3 configurations with separations between about 12 Å and 15 Å.
Figure 1: Comparison of the short-range (SR) and long-range (LR) machine learning interatomic potential performance for six dimer classes: charged-charged (CC), charged-polar (CP), polar-polar (PP), charged-apolar (CA), polar-apolar (PA), apolar-apolar (AA). For each class, the upper panel shows a snapshot of the system with the charge states indicated, the middle panel shows the parity plot for the force components, and the lower panel shows the binding energy curve, which is the potential energy difference between the dimer and two isolated monomers. The root mean square errors (RMSE) for the energy and force components of the test sets are shown in the insets.
For each molecular pair, we trained a short-range (SR) model using CACE with a cutoff $r_{\mathrm{cut}}=5$, 6 Bessel radial functions, $c=8$, $l_{\mathrm{max}}=2$, $\nu_{\mathrm{max}}=2$, $N_{\mathrm{embedding}}=3$, and one message passing layer ($T=1$). It is worth noting that this “SR” model setting already achieves a perceptive field of 10 Å via the message passing layer, which is quite typical for current MPNNs. In comparison, more traditional MLIPs based on local atomic descriptors typically use a cutoff of around 5 or 6 Å, making them even more short-ranged.
In the long-range (LR) model, the short-range component $E^{sr}$ used the same CACE setup, while the long-range component $E^{lr}$ employed a 4-dimensional hidden variable computed from the same CACE $B$-features and utilized Ewald summation (Eqn. (4)) with $\sigma=1$ and a $\mathbf{k}$-point cutoff of $k_{c}=2\pi/3$. It is important to fit both energy and forces for these datasets, as fitting only to a few energy values may result in models that accurately predict binding energy but perform poorly on forces.
Fig. 1 demonstrates that the LR MLIPs outperform the SR MLIPs across all cases. The improvement is particularly pronounced for the CC, CP, and PP cases, where the root mean square errors (RMSEs) of energy and force components can be reduced by an order of magnitude. Additionally, the SR models fail to adequately capture the training data, as indicated by the flattening of the binding curves and the large discrepancies between some predicted and true forces (gray symbols in Fig. 1). The primary limitation of the SR models is that molecules separated by distances beyond the MLIP cutoff exist on independent atomic graphs, rendering the message passing layers ineffective for communication between the dimers. The improvement achieved by the LR models for the CA, PA, and AA cases is more modest, though the energy scales in these systems are much smaller, and the absolute RMSE values are relatively low. The residual RMSEs can also be attributed to the limited (only 10) training data points used in each case. Direct comparison of the accuracy of the present LR models with the LODE method Huguenin-Dumittan et al. (2023) is challenging, as different training protocols were employed in Ref. Huguenin-Dumittan et al. (2023), and LODE depends significantly on the choice of the potential exponent $p$ for each dimer class. However, based solely on RMSE values, the LR models presented here appear to be more accurate. For example, in the CC class, the energy RMSE is 15.5 meV, compared to approximately 0.1 eV for LODE with the optimal choice of $p$.
III.2 Bulk water
As an example application to dipolar fluids, we applied the LES method to a dataset of 1,593 liquid water configurations, each containing 64 molecules Cheng et al. (2019). The dataset was computed using revPBE0-D3 DFT. For the CACE representation, we employed a cutoff of $r_{\mathrm{cut}}=5.5$, 6 Bessel radial functions with $c=12$, $l_{\mathrm{max}}=3$, $\nu_{\mathrm{max}}=3$, $N_{\mathrm{embedding}}=3$, and either no message passing ($T=0$) or one message passing layer ($T=1$). For the long-range component, we used a 4-dimensional $q$, a maximum cutoff of $k_{c}=\pi$, and $\sigma=1$ in the Ewald summation.
Figure 2: Learning curves of energy ($E$) and force ($F$) mean absolute errors (MAEs) on the bulk water dataset Cheng et al. (2019), using short-range (SR) or long-range (LR) models and with or without massage passing layers ($T=0$ or $T=1$).
The learning curves in Fig. 2 demonstrate that message passing significantly improves the accuracy of the MLIPs, consistent with previous studies Batzner et al. (2022); Batatia et al. (2022). The LR component further reduces the error for both models without a message passing layer ($T=0$) and with one message passing layer ($T=1$). The improvement is particularly notable in the $T=0$ scenario. For the $T=1$ models, the LR component results in a smaller reduction in errors, likely because the $T=1$ SR models already capture atomic interactions up to 11 Å. Nevertheless, the $T=1$ LR model demonstrates greater efficiency in learning with fewer data. We also compared the inference speeds of these models during molecular dynamics (MD) simulations, as detailed in the Methods section.
Figure 3: Predicted oxygen-oxygen radial distribution functions (RDF) of water at 300 K and 1 g/mL, using short-range (SR) or long-range (LR) models and with or without massage passing layers ($T=0$ or $T=1$). The experimental O-O RDF at ambient conditions was obtained from Ref Skinner et al. (2014).
To investigate how message passing and long-range interactions affect predicted structural properties, we performed MD simulations of bulk water with a density of 1 g/mL at 300 K using each model. The upper panel of Fig. 3 shows the oxygen-oxygen (O-O) radial distribution function (RDF) computed with different MLIPs. All the computed O-O RDFs are indistinguishable and in excellent agreement with experimental results from X-ray diffraction measurements Skinner et al. (2014). This suggests that local representations are sufficient for accurately predicting the RDFs of bulk liquid water, consistent with previous findings Yue et al. (2021).
Figure 4: Predicted dipole density correlations in reciprocal space, using short-range (SR) or long-range (LR) models and with or without massage passing layers ($T=0$ or $T=1$). The inset shows a zoom-in of the small $k$ values.
We then investigated the effect of long-range interactions on the dielectric properties of water by computing the longitudinal component of the dipole density correlation function Cox (2020), $\langle\tilde{m}_{z}^{\star}(k)\tilde{m}_{z}(k)\rangle$, where $\tilde{m}$ is the Fourier transform of the molecular dipole density and $\mathbf{k}=k\mathbf{z}$ is along the z-axis. The dipole correlation function at the long-wavelength limit can be used to determine the dielectric constant of water Gao and Remsing (2022).
As shown in Fig. 4, while the dipole density correlation functions predicted by different MLIPs are in excellent agreement for most values of $k$, discrepancies emerge at long wavelengths. Specifically, the results from short-ranged models sharply increase as $k\rightarrow 0$. This divergence has also been observed in previous studies comparing SR and LR models Cox (2020); Gao and Remsing (2022). The divergence of the $T=0$ short-ranged MLIP appears at a $k$ value corresponding to a real length scale of about 11 Å, while the $T=1$ SR model diverges at about 16 Å. Interestingly, these length scales, where divergence occurs, exceed the effective cutoff radius, suggesting that SR MLIPs may partially describe long-range effects beyond their cutoffs in a mean-field manner. Comparing the two SR MLIPs, the message passing layer delays the onset of divergence at small $k$ values but does not eliminate it. In this scenario, only true LR models can adequately describe the dielectric response.
Figure 5: Learning curves on the liquid–vapor water interface dataset Niblett et al. (2021) using the short-range (SR) and long-range (LR) models with no message passing ($T=0$) or one message passing layers ($T=1$).
III.3 Interfacial water
It is widely recognized that short-range MLIPs fall short in describing interfaces Niblett et al. (2021); Gao and Remsing (2022). To demonstrate the efficacy of the LES method for interfaces, we used a liquid-water interface dataset from Ref. Niblett et al. (2021), computed with the revPBE-D3 functional. We selected a subset of 500 liquid-vapor interface configurations, each containing 522 water molecules, far fewer than the approximately 17,500 training configurations used in Ref. Niblett et al. (2021) to train the DeePMD model Zhang et al. (2022), which may be attributed to the learning efficiency of CACE.
We employed the same settings for fitting the MLIPs as used for the bulk water dataset. The learning curves for forces are shown in Fig. 5. The energy errors are very low, reaching less than 0.1 meV/atom in mean absolute error (MAE) for all models using just 50 training configurations. The force errors are also small, ranging from about 14 meV/Å to 27 meV/Å in MAE when trained on 90% of the dataset. The learning behavior is similar to that observed in the bulk water case: both message passing and long-range interactions enhance the accuracy of the MLIPs.
Figure 6: A comparison of the interface properties of water predicted using short-range (SR) or long-range (LR) models and with or without massage passing layers ($T=0$ or $T=1$). a shows a snapshot of the thinner water slab configuration, b shows the water density profile, and c shows average cosine (c) the average $\cos(\theta)$ for the angle formed by the water dipole moment and z-axis. DFT results are from Ref. Niblett et al. (2021). d, e and f show the snapshot, water density, and $\langle\cos(\theta)\rangle$ for the thicker water slab, respectively.
For each of the $T=0$ SR, $T=0$ LR, $T=1$ SR, and $T=1$ LR models, we trained three MLIPs using different random splits of 90% for training and 10% for testing. We then simulated the liquid–vapor interface at 300 K using both a thinner slab (Fig. 6a) and a thicker slab of about double the thickness (Fig. 6d). The resulting density profiles of the water slabs are shown in Fig. 6b,e. All SR and LR models predict similar density profiles, and for the thinner slab, all models accurately reproduce the density profiles of the reference DFT water data.
In addition to density profiles, we evaluated the orientational order profiles, shown in Fig. 6c,f, by computing the angle, $\theta$, between the dipole orientation of water molecules and the $z$-axis. For each model, the three different fits of MLIPs provide slightly different predictions, and the variances are larger for the SR models. At the interfaces, water molecules tend to form a dipole layer, which is screened by subsequent layers, ensuring that the bulk does not exhibit net dipole moments Niblett et al. (2021). However, without long-range interactions, this screening effect is not adequately captured, leading to extended dipole ordering into the bulk as seen in the $T=0$ SR model. The introduction of message passing ($T=1$ SR model) alleviates this issue at least for the thinner slab, but it also introduces an artifact: dipole ordering in the bulk along the opposite direction for the thicker slab. In contrast, the LR models recover the correct orientational order profiles, effectively capturing the screening effect and accurately representing the physical polarization behavior at the liquid-vapor interfaces.
IV Discussion and Conclusions
In summary, we present a simple and general method, the Latent Ewald Summation (LES), for incorporating long-range interactions in atomistic systems. Unlike existing LR methods, LES does not require user-defined electrostatics or dispersion baseline corrections Niblett et al. (2021); Lee et al. (2022); Unke et al. (2021b), does not rely on partial charges or Wannier centers during training Unke and Meuwly (2019); Ko et al. (2021); Gao and Remsing (2022); Sifain et al. (2018); Gong et al. (2024); Shaidu et al. (2024), and does not utilize charge equilibration schemes Rappe and Goddard III (1991). We hypothesize that there may be underlying mathematical connections between the LES method and other approaches like LODE Grisafi and Ceriotti (2019); Huguenin-Dumittan et al. (2023) or recent density-based long-range electrostatic descriptors Faller et al. (2024), though LES offers a simpler mathematical formulation and implementation.
Our results demonstrate that the LR model effectively reproduces correct dimer binding curves for pairs of charged, polar, and apolar molecules, long-range dipole correlations in liquid water, and dielectric screening at liquid-vapor interfaces. In contrast, standard short-ranged MLIPs, even when utilizing message passing, often yield unphysical predictions in these contexts. These challenges are common in atomistic simulations of molecules and materials, particularly in systems like water, which is ubiquitous in nature and essential to biological processes. Moreover, long-range effects are crucial in aqueous solutions, organic electrolytes, and other interfacial systems.
We anticipate that the LES method will be widely adopted in atomistic modeling, especially for systems with significant electrostatic and dielectric properties. The long-range MLIP incurs only a modest computational cost, roughly double that of the short-range version (see Methods). Additionally, the LES method can be easily integrated into other MLIP frameworks, such as HDNNP Behler and Parrinello (2007), ACE Drautz (2019), GAP Bartók et al. (2010), MTPs Shapeev (2016), DeePMD Zhang et al. (2022), NequIP Batzner et al. (2022), MACE Batatia et al. (2022), etc.
V Methods
Implementation
We implemented the LES method as an Ewald module in CACE, which is written using PyTorch. The code is available in https://github.com/BingqingCheng/cace. The current implementation of the Ewald summation should in principle follow an overall scaling of $\mathcal{O}(N^{3/2})$ Toukmaji and Board Jr (1996). The implementation of the code can be further optimized by using algorithms with $\mathcal{O}(N\log(N))$ scaling Toukmaji and Board Jr (1996). The MD simulations can be performed in the ASE package, but the LAMMPS interface implementation is currently absent.
Benchmark of inference speed
We benchmarked short-range and long-range CACE water models with and without message passing for MD simulations of liquid water using a single Nvidia L40S GPU with 48 GB of memory. Fig. 7 illustrates the time required per MD step. All models, with and without LR or message passing, exhibit near-linear scaling. SR models support simulations with up to approximately 30,000 atoms on a single GPU, while LR models handle around 10,000 atoms due to the higher memory demands of the Ewald routine, which may still be optimized. LR models run at roughly half the speed of SR models. It may be sufficient to use a lower $\mathbf{k}$-point cutoff in the LR model to further increase the speed.
Figure 7: Timing of molecular dynamics simulations of the bulk liquid water system with varying number of atoms using different CACE models on a L40S GPU.
MD of Bulk Water
For each MLIP, we performed NVT simulations of bulk water at 1 g/mL and 300 K to compute the RDF shown in Fig. 3. The simulation cell contained 512 water molecules, with a time step of 1 femtosecond, using the Nosé-Hoover thermostat. The total simulation time was 300 ps.
To compute the longitudinal component of the dipole density correlation function for each model, we conducted an NVT simulation of 2,048 water molecules at 1 g/mL and 300 K in an elongated simulation box with a z-dimension of 99.3 Å. The simulation time for this system was 200 ps.
To compute the dipole density correlation functions, we calculated the dipole moment of each water molecule by assuming that the hydrogen atoms and oxygen atoms have charges of $+0.4238e$ and $-0.8476e$, respectively. This assumption only affects the absolute amplitude of the correlation function in Fig. 4, without altering the relative scale.
MD of Water Liquid-Vapor Interfaces
For the NVT simulations of the thinner slab at 300 K, we used a simulation cell with dimensions $(25.6~{}\mathrm{\AA},25.6~{}\mathrm{\AA},65.0~{}\mathrm{\AA})$, containing 522 water molecules. The thicker slab system had dimensions $(24.8~{}\mathrm{\AA},24.8~{}\mathrm{\AA},120.0~{}\mathrm{\AA})$, containing 1,024 water molecules. In both cases, the width of the water slab was less than half the total width of the cell. For the thinner slab, each independent NVT simulation lasted for about 400 ps. One run using $T=1$ SR model became unstable after about 280 ps and we discarded the unstable portion. For thicker slab, each run lasted for about 600 ps. One simulation using $T=1$ SR model became unstable after about 250 ps, and we only used the stable portion for the analysis.
Data availability
The training scripts, trained CACE potentials, and MD input files are available at https://github.com/BingqingCheng/cace-lr-fit.
Code availability The CACE package is publicly available at https://github.com/BingqingCheng/cace.
Competing Interests BC has an equity stake in AIMATX Inc.

### v2

@@
 Machine learning interatomic potentials (MLIPs) often neglect long-range interactions, such as electrostatic and dispersion forces.
-In this work, we introduce a straightforward and efficient method to account for long-range interactions by learning a latent variable from local atomic descriptors and applying an Ewald summation to this variable.
-We demonstrate that in systems including charged, polar, or apolar molecular dimers, bulk water, and water-vapor interface, standard short-ranged MLIPs can lead to unphysical predictions even when employing message passing.
+In this work, we introduce a straightforward and efficient method to account for long-range interactions by learning a hidden variable from local atomic descriptors and applying an Ewald summation to this variable.
+We demonstrate that in systems including charged and polar molecular dimers, bulk water, and water-vapor interface, standard short-ranged MLIPs can lead to unphysical predictions even when employing message passing.
 The long-range models effectively eliminate these artifacts, with only about twice the computational cost of short-range MLIPs.
-I Introduction
-Machine learning interatomic potentials (MLIPs) can learn from reference quantum mechanical calculations and then predict the energy and forces of atomic configurations quickly, thus allowing for a more accurate and comprehensive exploration of material and molecular properties at scale Keith et al.
+Machine learning interatomic potentials (MLIPs) can learn from reference quantum mechanical calculations and then predict the energy and forces of atomic configurations quickly, allowing for a more accurate and comprehensive exploration of material and molecular properties at scale Keith et al.
 (2021); Unke et al.
@@
 The short-range MLIPs, however, neglect all kinds of long-range interactions such as Coulomb and dispersion.
-Although short-range potentials may be sufficient for describing most properties of homogeneous bulk systems Yue et al.
+Although short-range potentials may be sufficient to describe most properties of homogeneous bulk systems Yue et al.
 (2021), they may fail for liquid-vapor interfaces Niblett et al.
@@
 (2021) predicts the electronegativities of each nucleus, and then use a charge equilibration scheme Rappe and Goddard III (1991) to assign the charges.
-4G-HDNNPs are trained directly to reproduce atomic partial charges from reference quantum mechanical calculations, although partial charges are not physical observables and their values depend on the specific partitioning scheme used Sifain et al.
+4G-HDNNPs are trained directly to reproduce atomic partial charges from reference quantum mechanical calculations, although partial charges are not physical observable and their values depend on the specific partitioning scheme used Sifain et al.
 (2018).
@@
 (2023); Haghighatlari et al.
-(2022) employ a number of graph convolution layers to communicate information between atoms, thereby capturing long-range interaction up to the local cutoff radius times the number of layers.
+(2022) employ a number of graph convolution layers to communicate information between atoms, thus capturing long-range interaction up to the local cutoff radius times the number of layers.
 However, if parts of the system are disconnected on the graph, e.g.
 two molecules with a distance beyond the cutoff, the message passing scheme does not help.
-A very interesting approach is the long-distance equivariant (LODE) method Grisafi and Ceriotti (2019); Huguenin-Dumittan et al.
-(2023), which uses local descriptors to encode the Coulomb and other asymptotic decaying potentials ($1/r^{p}$) around the atoms.
+Another class of methods is to learn the long-range descriptors and interactions in the reciprocal space with learnable frequency filters Yu et al.
+(2022); Kosmala et al.
+(2023).
+Finally, a very interesting approach is the long-distance equivariant (LODE) method Grisafi and Ceriotti (2019); Huguenin-Dumittan et al.
+(2023), which uses local descriptors to encode the Coulomb and other asymptotic decaying potentials ($1/r^{p}$) around the atoms, and a related, density-based long-range descriptors Faller et al.
+(2024).
 Here, we propose a simple method, the Latent Ewald Summation (LES), for accounting for long-range interactions of atomistic systems.
@@
 After describing the algorithm, we benchmark LES on selected molecular and material systems.
-II Theory
+I Results
+I.1 Theory
 For a periodic atomic system, the total potential energy is decomposed into a short-range and a long-range part, i.e.
@@
 where $E_{\theta}$ is a multilayer perceptron with parameters $\theta$ that maps the invariant features ($B$) of an atom to its short-range atomic energy.
-$B$ can be any invariant features used in different MLIP methods, including those based on local atomic environment descriptors such as ACE Drautz (2019), atom centered symmetry functions Behler and Parrinello (2007), smooth overlap of atomic positions (SOAP) Bartók et al.
+$B$ can be any invariant features used in different MLIP methods, including those based on local atomic environment descriptors such as ACE Drautz (2019), atom-centered symmetry functions Behler and Parrinello (2007), smooth overlap of atomic positions (SOAP) Bartók et al.
 (2013), or any latent invariant features in MPNNs.
@@
 where $\mathbf{k}=(2\pi n_{x}/L_{x},2\pi n_{y}/L_{y},2\pi n_{z}/L_{z})$ is a reciprocal vector of the orthorhombic cell, and $\mathbf{r}_{i}$ is the Cartesian coordinates of atom $i$.
-The long-range energy is then obtained using the Ewald summation:
+The long-range energy is then obtained using an Ewald summation form that best captures the electrostatic potential ($1/r$Williams (1971) :
 $E^{lr}=\dfrac{1}{V}\sum_{0<k<k_{c}}\dfrac{e^{-\sigma^{2}k^{2}/2}}{k^{2}}|S(\mathbf{k})|^{2},$ (4)
-where $\sigma$ is a smearing factor which we typically set to 1 Å, $k=|\mathbf{k}|$ is the magnitude, and $k_{c}$ is the maximum cutoff.
+where $\sigma$ is a smearing factor which we typically set to 1 Å with justifications in the Methods , $k=|\mathbf{k}|$ is the magnitude, and $k_{c}$ is the maximum cutoff.
 $q$ can be multi-dimensional, in which case the total long-range energy is aggregated over contributions from different dimensions of $q$ after the Ewald summation.
 The LES method can be interpreted in two ways.
-First, the hidden variable $q$ is analogous to environment-dependent partial charges on each atom.
+First, the hidden variable $q$ is analogous to the environmental-dependent partial charges on each atom.
 This implies that the method is at least as expressive as those explicitly based on learning partial atomic charges, as it would yield the same results if $q$ replicated the partial charges.
-Additionally, $q$ can be multi-dimensional, potentially enhancing expressiveness further.
+Additionally, $q$ can be multidimensional, potentially enhancing expressiveness further.
 Unlike partial charges, $q$ is not constrained by requirements such as charge neutrality or correct charge magnitudes.
+As the Ewald summation in Eqn.
+(4) omits the $k=0$ term, a non-zero net $q$ does not cause energy divergence issues.
 Since the Ewald summation of $q$ doesn’t need to correspond to a physical electrostatic potential, Eqn.
@@
 The second interpretation of LES is as a mechanism that allows atoms far apart in the simulation box to communicate their local information.
-In this sense, LES is related to the recent Ewald-based long-range message-passing method, which facilitates message exchange between atoms in reciprocal space Kosmala et al.
+In this sense, LES is related to the recent Ewald-based long-range message passing method, which facilitates message exchange between atoms in reciprocal space Kosmala et al.
 (2023).
-III Examples
-III.1 Molecular dimers
-We benchmark the LES method on the binding curves between dimers of charged (C), polar (P), and apolar (A) relaxed molecules at various separations in a periodic cubic box with a 30 Å edge length.
+I.2 Example on molecular dimers
+We benchmark the LES method on the binding curves between dimers of charged (C) and polar (P) molecules at various separations in a periodic cubic box with a 30 Å edge length.
 The dataset Huguenin-Dumittan et al.
-(2023) includes energy and force information calculated using the HSE06 hybrid density functional theory (DFT) with a many-body dispersion correction.
-We selected one example from each of the six dimer classes (CC, CP, PP, CA, CP, AA), derived from the combination of the three monomer categories.
+(2023), originally from the BioFragment Database Burns et al.
+(2017), includes energy and force information calculated using the HSE06 hybrid density functional theory (DFT) with a many-body dispersion correction.
+We selected one example from each of the three dimer classes (CC, CP, PP) , derived from the combination of the three monomer categories.
 Figure 1 shows a snapshot of each example.
 The goal of this benchmark is to evaluate whether the MLIP models can extrapolate dimer interactions at larger separations based on training data from smaller distances.
-For each molecular pair, the training set consists of 10 configurations with dimer separation distances between approximately 5 Å and 12 Å, and the testing set includes 3 configurations with separations between about 12 Å and 15 Å.
-Figure 1: Comparison of the short-range (SR) and long-range (LR) machine learning interatomic potential performance for six dimer classes: charged-charged (CC), charged-polar (CP), polar-polar (PP), charged-apolar (CA), polar-apolar (PA), apolar-apolar (AA).
-For each class, the upper panel shows a snapshot of the system with the charge states indicated, the middle panel shows the parity plot for the force components, and the lower panel shows the binding energy curve, which is the potential energy difference between the dimer and two isolated monomers.
+For each molecular pair, the training set consists of 10 configurations with dimer separation distances between approximately 5 Å and 12 Å, and the test set includes 3 configurations with separations between approximately 12 Å and 15 Å.
+Figure 1: Comparison of the short-range (SR) and long-range (LR) machine learning interatomic potential performance for three dimer classes: charged-charged (CC), charged-polar (CP), and polar-polar (PP).
+For each class, the upper panel shows a snapshot of the system with the charge states indicated, the middle panel shows the parity plot for the force components, and the lower panel shows the binding energy curve, which is the potential energy difference between the dimer, and two isolated and relaxed monomers.
 The root mean square errors (RMSE) for the energy and force components of the test sets are shown in the insets.
 For each molecular pair, we trained a short-range (SR) model using CACE with a cutoff $r_{\mathrm{cut}}=5$, 6 Bessel radial functions, $c=8$, $l_{\mathrm{max}}=2$, $\nu_{\mathrm{max}}=2$, $N_{\mathrm{embedding}}=3$, and one message passing layer ($T=1$).
-It is worth noting that this “SR” model setting already achieves a perceptive field of 10 Å via the message passing layer, which is quite typical for current MPNNs.
+It should be noted that this setting of the “SR” model already achieves a perceptive field of 10 Å through the message passing layer, which is quite typical for current MPNNs.
 In comparison, more traditional MLIPs based on local atomic descriptors typically use a cutoff of around 5 or 6 Å, making them even more short-ranged.
@@
 Fig.
-1 demonstrates that the LR MLIPs outperform the SR MLIPs across all cases.
-The improvement is particularly pronounced for the CC, CP, and PP cases, where the root mean square errors (RMSEs) of energy and force components can be reduced by an order of magnitude.
-Additionally, the SR models fail to adequately capture the training data, as indicated by the flattening of the binding curves and the large discrepancies between some predicted and true forces (gray symbols in Fig.
+1 shows that the LR MLIPs outperform the SR MLIPs in all cases.
+Furthermore, SR models fail to adequately capture the training data, as indicated by flattening of the binding curves and the large discrepancies between some predicted and true forces (gray symbols in Fig.
 1).
-The primary limitation of the SR models is that molecules separated by distances beyond the MLIP cutoff exist on independent atomic graphs, rendering the message passing layers ineffective for communication between the dimers.
-The improvement achieved by the LR models for the CA, PA, and AA cases is more modest, though the energy scales in these systems are much smaller, and the absolute RMSE values are relatively low.
-The residual RMSEs can also be attributed to the limited (only 10) training data points used in each case.
+The primary limitation of the SR models is that molecules separated by distances beyond the MLIP cutoff exist on independent atomic graphs, rendering the message-passing layers ineffective for communication between the dimers.
 Direct comparison of the accuracy of the present LR models with the LODE method Huguenin-Dumittan et al.
-(2023) is challenging, as different training protocols were employed in Ref.
+(2023) is challenging, as different training protocols were used in Ref.
 Huguenin-Dumittan et al.
@@
 For example, in the CC class, the energy RMSE is 15.5 meV, compared to approximately 0.1 eV for LODE with the optimal choice of $p$.
-III.2 Bulk water
+I.3 Example on molten NaCl
+Molten bulk sodium chloride (NaCl) presents non-negligible long-range electrostatic interactions.
+The dataset from Ref.
+Faller et al.
+(2024) contains 1014 structures (80% train and 20% validation) of 64 Na and 64 Cl atoms.
+Table 1 compares the root-mean-square percentage error (RMSPE) using different methods.
+$r_{\mathrm{cut}}=~{}$6 Å were used for all models except for the LODE flexible (7 Å).
+The CACE-LR models use 6 Bessel radial functions, $c=12$, $l_{\mathrm{max}}=3$, $\nu_{\mathrm{max}}=3$, $N_{\mathrm{embedding}}=3$, zero ($T=0$) or one message passing layer ($T=1$), a 4-dimensional $q$, $\sigma=1$ and a $\mathbf{k}$-point cutoff of $k_{c}=2\pi/3$ in Ewald summation.
+The details for the other models are in Ref.
+Faller et al.
+(2024).
+Table 1: The root-mean-square percentage error (RMSPE) for the validation set of molten NaCl, using the short-ranged SOAP descriptor Bartók et al.
+(2010, 2013); Jinnouchi et al.
+(2019), two setting of LODE Grisafi and Ceriotti (2019); Huguenin-Dumittan et al.
+(2023) descriptors (minimal and flexible), MACE Batatia et al.
+(2022) with zero ($T=0$) or one message passing layers ($T=1$) as well as using invariant and equivariant messages, a density-based long-range model (Density-LR) Faller et al.
+(2024), and CACE-LR with zero ($T=0$) or one message passing layers ($T=1$).
+The fits other than the CACE models are from Ref.
+Faller et al.
+(2024).
+RMSPE E (%) RMSPE F (%)
+SOAP Bartók et al. (2010, 2013); Jinnouchi et al. (2019) 7.7 12.6
+MACE Batatia et al. (2022) T=0 8.5 8.5
+MACE Batatia et al. (2022) T=1 invariant 3.0 3.1
+MACE Batatia et al. (2022) T=1 equivariant 2.1 2.2
+LODE Grisafi and Ceriotti (2019); Huguenin-Dumittan et al. (2023) minimal 5.6 3.7
+LODE Grisafi and Ceriotti (2019); Huguenin-Dumittan et al. (2023) flexible 7.0 11.0
+Density-LR Faller et al. (2024) 2.2 3.3
+CACE-LR T=0 1.9 2.3
+CACE-LR T=1 1.4 1.8
+Table 1 shows that the short-ranged models, including SOAP Bartók et al.
+(2010, 2013); Jinnouchi et al.
+(2019) and MACE Batatia et al.
+(2022) $T=0$, have the worst performances.
+All the long-ranged models, including LODE Grisafi and Ceriotti (2019); Huguenin-Dumittan et al.
+(2023) with different settings, a density-based long-range model Faller et al.
+(2024), and CACE-LR outperform the purely SR methods.
+MACE Batatia et al.
+(2022) with message passing ($T=1$), especially the one using equivariant features, also achieves accurate results, as message passing increases the effective perceptive field.
+Overall, CACE-LR models, with and without message passing ($T=0$ and $T=1$), obtain the best accuracy.
+I.4 Example on bulk water
 As an example application to dipolar fluids, we applied the LES method to a dataset of 1,593 liquid water configurations, each containing 64 molecules Cheng et al.
 (2019).
-The dataset was computed using revPBE0-D3 DFT.
-For the CACE representation, we employed a cutoff of $r_{\mathrm{cut}}=5.5$, 6 Bessel radial functions with $c=12$, $l_{\mathrm{max}}=3$, $\nu_{\mathrm{max}}=3$, $N_{\mathrm{embedding}}=3$, and either no message passing ($T=0$) or one message passing layer ($T=1$).
+The dataset was calculated using revPBE0-D3 DFT.
+For the CACE representation, we used a cutoff of $r_{\mathrm{cut}}=5.5$, 6 Bessel radial functions with $c=12$, $l_{\mathrm{max}}=3$, $\nu_{\mathrm{max}}=3$, $N_{\mathrm{embedding}}=3$, and no message passing ($T=0$) or no message passing layer ($T=1$).
 For the long-range component, we used a 4-dimensional $q$, a maximum cutoff of $k_{c}=\pi$, and $\sigma=1$ in the Ewald summation.
@@
 The learning curves in Fig.
-2 demonstrate that message passing significantly improves the accuracy of the MLIPs, consistent with previous studies Batzner et al.
+2 demonstrate that message passing significantly improves the accuracy of MLIPs, consistent with previous studies Batzner et al.
 (2022); Batatia et al.
 (2022).
-The LR component further reduces the error for both models without a message passing layer ($T=0$) and with one message passing layer ($T=1$).
+The LR component further reduces the error for both models without a message passing layer ($T=0$) and with a message passing layer ($T=1$).
 The improvement is particularly notable in the $T=0$ scenario.
-For the $T=1$ models, the LR component results in a smaller reduction in errors, likely because the $T=1$ SR models already capture atomic interactions up to 11 Å.
+For the $T=1$ models, the LR component results in a smaller reduction in errors, probably because the $T=1$ SR models already capture atomic interactions up to 11 Å.
 Nevertheless, the $T=1$ LR model demonstrates greater efficiency in learning with fewer data.
@@
 3 shows the oxygen-oxygen (O-O) radial distribution function (RDF) computed with different MLIPs.
-All the computed O-O RDFs are indistinguishable and in excellent agreement with experimental results from X-ray diffraction measurements Skinner et al.
+All computed O-O RDFs are indistinguishable and in excellent agreement with the experimental results of the X-ray diffraction measurements Skinner et al.
 (2014).
-This suggests that local representations are sufficient for accurately predicting the RDFs of bulk liquid water, consistent with previous findings Yue et al.
+This suggests that local representations are sufficient to accurately predict the RDFs of bulk liquid water, consistent with previous findings Yue et al.
 (2021).
@@
 Specifically, the results from short-ranged models sharply increase as $k\rightarrow 0$.
-This divergence has also been observed in previous studies comparing SR and LR models Cox (2020); Gao and Remsing (2022).
-The divergence of the $T=0$ short-ranged MLIP appears at a $k$ value corresponding to a real length scale of about 11 Å, while the $T=1$ SR model diverges at about 16 Å.
+This divergence has also been observed in previous studies comparing the SR and LR models Cox (2020); Gao and Remsing (2022).
+The divergence of the $T=0$ short-ranged MLIP appears at a $k$ value corresponding to a real length scale of approximately 11 Å, while the $T=1$ SR model diverges at about 16 Å.
 Interestingly, these length scales, where divergence occurs, exceed the effective cutoff radius, suggesting that SR MLIPs may partially describe long-range effects beyond their cutoffs in a mean-field manner.
-Comparing the two SR MLIPs, the message passing layer delays the onset of divergence at small $k$ values but does not eliminate it.
+When comparing the two SR MLIPs, the message passing layer delays the onset of divergence at small $k$ values but does not eliminate it.
 In this scenario, only true LR models can adequately describe the dielectric response.
@@
 (2021) using the short-range (SR) and long-range (LR) models with no message passing ($T=0$) or one message passing layers ($T=1$).
-III.3 Interfacial water
+I.5 Example on interfacial water
 It is widely recognized that short-range MLIPs fall short in describing interfaces Niblett et al.
@@
 5.
-The energy errors are very low, reaching less than 0.1 meV/atom in mean absolute error (MAE) for all models using just 50 training configurations.
+The energy errors are very low, reaching less than 0.1 meV/atom in mean absolute error (MAE) for all models using only 50 training configurations.
 The force errors are also small, ranging from about 14 meV/Å to 27 meV/Å in MAE when trained on 90% of the dataset.
@@
 (2021).
-d, e and f show the snapshot, water density, and $\langle\cos(\theta)\rangle$ for the thicker water slab, respectively.
+d, e and f show the snapshot, the water density, and $\langle\cos(\theta)\rangle$ for the thicker water slab, respectively.
 For each of the $T=0$ SR, $T=0$ LR, $T=1$ SR, and $T=1$ LR models, we trained three MLIPs using different random splits of 90% for training and 10% for testing.
@@
 6c,f, by computing the angle, $\theta$, between the dipole orientation of water molecules and the $z$-axis.
-For each model, the three different fits of MLIPs provide slightly different predictions, and the variances are larger for the SR models.
+For each model, the three different fits of MLIPs provide slightly different predictions, and the variances are larger for the SR models compared to the LR models.
 At the interfaces, water molecules tend to form a dipole layer, which is screened by subsequent layers, ensuring that the bulk does not exhibit net dipole moments Niblett et al.
@@
 The introduction of message passing ($T=1$ SR model) alleviates this issue at least for the thinner slab, but it also introduces an artifact: dipole ordering in the bulk along the opposite direction for the thicker slab.
-In contrast, the LR models recover the correct orientational order profiles, effectively capturing the screening effect and accurately representing the physical polarization behavior at the liquid-vapor interfaces.
-IV Discussion and Conclusions
+In contrast, on average and with smaller variance between different fits , the LR models recover the correct orientational order profiles, effectively capturing the screening effect and accurately representing the physical polarization behavior at the liquid-vapor interfaces.
+II Discussion
+The current LES method can be extended in a number of ways.
+The form of the Ewald summation in Eqn.
+(4) was selected to best capture the electrostatic potential ($1/r$), which dominates long-range interactions.
+It remains unclear what is the effect of not directly accounting for other decaying potentials.
+Different forms can be used to capture other decaying potentials $1/r^{p}$ with different $p$ components.
+For examples, to best capture the London dispersion ($1/r^{6}$), one can use Chen et al.
+(1997); Williams (1971).
+$E_{6}^{lr}=\left(\frac{\pi^{3/2}}{24V}\right)\sum_{k>0}k^{3}\\ \left[\pi^{1/2}\mathrm{erfc}(b)+\left(\frac{1}{2b^{3}}-\frac{1}{b}\right)\times\exp(-b^{2})\right]|S(\mathbf{k})|^{2},$ (5)
+where $b^{2}=\sigma^{2}k^{2}/2$, and $\mathrm{erfc}$ denotes the complimentary error function.
+Currently, the hidden variable $q$ is rotational invariant, and it may be extended to adopt a vector or tensor form to capture the dipole and multiple moments on an atom.
+The relation between LES and the LODE method Grisafi and Ceriotti (2019); Huguenin-Dumittan et al.
+(2023) is worth exploring.
+The key difference is the ordering of local and global interactions.
+In LES, the hidden variable $q_{i}$ is determined by the local atomic environment-dependent features $B_{i}$ of atom $i$, reflecting the nearsightedness principle: partial atomic charges depend mainly on local environments, similar to the short-range approximation in standard MLIPs.
+Although it should be pointed out that such locality approximation remains empirical rather than systematic.
+These local $q$ then interact globally via the Ewald summation to determine the global long-range energy.
+In contrast, in LODE Grisafi and Ceriotti (2019); Huguenin-Dumittan et al.
+(2023) the potential field (e.g.
+electrostatic potential and other $1/r^{p}$ decaying potentials) generated by all the atoms in the system is calculated in the reciprocal space via Ewald summation, and such field near a central atom $i$ up to some cutoff radius is then projected onto a set of basis functions to form the LODE descriptors.
+Finally, the short-range descriptors and the LODE descriptors of atom $i$ are used to predict its energy.
+The density-based long-range descriptors Faller et al.
+(2024) are similar to LODE, but the global atomic density itself is used instead of the field.
+In summary, LES uses local descriptors to predict global long-range energy, while LODE uses global field to inform local descriptors.
+There may be underlying mathematical connections between the LES and LODE, but this is not obvious, particularly considering the non-linearity introduced by the multilayer perceptron in LES that predicts $q_{i}$ from $B_{i}$ (Eqn.
+(2)).
+LES also shares similarities with the recently proposed Ewald-based long-range message-passing method Kosmala et al.
+(2023), as both utilize structure factors.
+While LES directly determines the long-range energy from the structure factor of a hidden low-dimensional variable (Eqn.(4)), the Ewald message-passing method Kosmala et al.
+(2023) employs a learnable frequency filter to map the structure factor of atomic descriptors to real space.
+This mapping generates a long-range message for each atom that is used to update atomic descriptors during the message passing step.
+LES can thus be interpreted as a minimalist version of Ewald message-passing, with only one long-range message passing layer, compressed message dimension, fixed frequency filter, and simplified readout.
+Such minimalist design makes the model light-weight, and allows for easier physical interpretation.
 In summary, we present a simple and general method, the Latent Ewald Summation (LES), for incorporating long-range interactions in atomistic systems.
-Unlike existing LR methods, LES does not require user-defined electrostatics or dispersion baseline corrections Niblett et al.
+Unlike many existing LR methods, LES does not require user-defined electrostatic or dispersion baseline corrections Niblett et al.
 (2021); Lee et al.
@@
 (2024), and does not utilize charge equilibration schemes Rappe and Goddard III (1991).
-We hypothesize that there may be underlying mathematical connections between the LES method and other approaches like LODE Grisafi and Ceriotti (2019); Huguenin-Dumittan et al.
-(2023) or recent density-based long-range electrostatic descriptors Faller et al.
-(2024), though LES offers a simpler mathematical formulation and implementation.
-Our results demonstrate that the LR model effectively reproduces correct dimer binding curves for pairs of charged, polar, and apolar molecules, long-range dipole correlations in liquid water, and dielectric screening at liquid-vapor interfaces.
+Moreover, LES shows best accuracy compared to other long-range method and message passing neutral networks in the example on molten salt.
+Our results demonstrate that the LR model effectively reproduces correct dimer binding curves for pairs of charged and polar molecules, long-range dipole correlations in liquid water, and dielectric screening at liquid-vapor interfaces.
 In contrast, standard short-ranged MLIPs, even when utilizing message passing, often yield unphysical predictions in these contexts.
@@
 The long-range MLIP incurs only a modest computational cost, roughly double that of the short-range version (see Methods).
-Additionally, the LES method can be easily integrated into other MLIP frameworks, such as HDNNP Behler and Parrinello (2007), ACE Drautz (2019), GAP Bartók et al.
+Furthermore, the LES method can be easily integrated into other MLIP frameworks, such as HDNNP Behler and Parrinello (2007), ACE Drautz (2019), GAP Bartók et al.
 (2010), MTPs Shapeev (2016), DeePMD Zhang et al.
@@
 (2022), etc.
-V Methods
-Implementation
+III Methods
+III.1 Implementation
 We implemented the LES method as an Ewald module in CACE, which is written using PyTorch.
 The code is available in https://github.com/BingqingCheng/cace.
-The current implementation of the Ewald summation should in principle follow an overall scaling of $\mathcal{O}(N^{3/2})$ Toukmaji and Board Jr (1996).
-The implementation of the code can be further optimized by using algorithms with $\mathcal{O}(N\log(N))$ scaling Toukmaji and Board Jr (1996).
+The current implementation of the Ewald summation should in principle follow an overall scaling of $\mathcal{O}(N^{3/2})$ Toukmaji and Board Jr (1996), with $N$ being the number of atoms.
+The implementation of the code can be further optimized using algorithms with $\mathcal{O}(N\log(N))$ scaling employing techniques with the charges interpolated to a 3D grid such as the Particle-Particle Particle-Mesh method  Toukmaji and Board Jr (1996).
 The MD simulations can be performed in the ASE package, but the LAMMPS interface implementation is currently absent.
-Benchmark of inference speed
-We benchmarked short-range and long-range CACE water models with and without message passing for MD simulations of liquid water using a single Nvidia L40S GPU with 48 GB of memory.
+III.2 Selection of the smearing factor $\sigma$
+For a system of point charges, the Ewald summation expression in Eqn.
+(4) implies that the long-range energy corresponds to the electrostatic energy of Gaussian distributed charges with $\sigma$ being the standard deviation, while the short-range term needs to account for the difference between the Gaussian charges and the point charges.
+The value of the smearing factor $\sigma$ needs to be a fraction of the real-space cutoff radius $r_{\mathrm{rut}}$ in order to effectively eliminate the short-range error Kolafa and Perram (1992).
+Meanwhile, a larger $\sigma$ makes the Ewald series in $\mathbf{k}$ (Eqn.
+(4)) more converged and avoids a high reciprocal space cutoff $k_{c}$.
+and its optimal choice has been extensively discussed in the literature Kolafa and Perram (1992).
+Combining these considerations, $\sigma$ values in the rough range of 0.5 Å and 2 Å are reasonable choices.
+We tested the choice of $\sigma$ on the bulk water dataset, with same settings as before and no message passing layer.
+Fig.
+7 shows that $\sigma=1$ in this case leads to the lowest error, although the differences in outcome between various $\sigma$ choices are quite small.
+Figure 7: Energy ($E$) and force ($F$) mean absolute errors (MAEs) on the bulk water dataset Cheng et al.
+(2019) with difference values of the smearing factor $\sigma$, using short-range models without massage passing.
+III.3 Benchmark of inference speed
+We benchmarked short- and long-range CACE water models with and without message passing for MD simulations of liquid water using a single Nvidia L40S GPU with 48 GB of memory.
 Fig.
 8 illustrates the time required per MD step.
-All models, with and without LR or message passing, exhibit near-linear scaling.
+All models, with and without LR or message passing, exhibit favorable scaling.
 SR models support simulations with up to approximately 30,000 atoms on a single GPU, while LR models handle around 10,000 atoms due to the higher memory demands of the Ewald routine, which may still be optimized.
@@
 It may be sufficient to use a lower $\mathbf{k}$-point cutoff in the LR model to further increase the speed.
-Figure 7: Timing of molecular dynamics simulations of the bulk liquid water system with varying number of atoms using different CACE models on a L40S GPU.
-MD of Bulk Water
+Figure 8: Timing of molecular dynamics simulations of the bulk liquid water system with varying number of atoms ($N$) on a L40S GPU, using short-range (SR) or long-range (LR) CACE models and with or without massage passing layers ($T=0$ or $T=1$).
+III.4 MD of Bulk Water
 For each MLIP, we performed NVT simulations of bulk water at 1 g/mL and 300 K to compute the RDF shown in Fig.
@@
 The total simulation time was 300 ps.
-To compute the longitudinal component of the dipole density correlation function for each model, we conducted an NVT simulation of 2,048 water molecules at 1 g/mL and 300 K in an elongated simulation box with a z-dimension of 99.3 Å.
+To compute the longitudinal component of the dipole density correlation function for each model, we performed an NVT simulation of 2,048 water molecules at 1 g/mL and 300 K in an elongated simulation box with a $z$-dimension of 99.3 Å.
 The simulation time for this system was 200 ps.
-To compute the dipole density correlation functions, we calculated the dipole moment of each water molecule by assuming that the hydrogen atoms and oxygen atoms have charges of $+0.4238e$ and $-0.8476e$, respectively.
+To compute the dipole density correlation functions, we calculated the dipole moment of each water molecule assuming that hydrogen atoms and oxygen atoms have charges of $+0.4238e$ and $-0.8476e$, respectively.
 This assumption only affects the absolute amplitude of the correlation function in Fig.
 4, without altering the relative scale.
-MD of Water Liquid-Vapor Interfaces
+III.5 MD of Water Liquid-Vapor Interfaces
 For the NVT simulations of the thinner slab at 300 K, we used a simulation cell with dimensions $(25.6~{}\mathrm{\AA},25.6~{}\mathrm{\AA},65.0~{}\mathrm{\AA})$, containing 522 water molecules.
@@
 One run using $T=1$ SR model became unstable after about 280 ps and we discarded the unstable portion.
-For thicker slab, each run lasted for about 600 ps.
-One simulation using $T=1$ SR model became unstable after about 250 ps, and we only used the stable portion for the analysis.
+For the thicker slab, each run lasted for about 600 ps.
+One simulation using the $T=1$ SR model became unstable after about 250 ps, and we only used the stable portion for the analysis.
 Data availability
@@
 Code availability The CACE package is publicly available at https://github.com/BingqingCheng/cace.
-Competing Interests BC has an equity stake in AIMATX Inc.
+Acknowledgements
+BC thanks David Limmer for providing the water slab dataset, and Carolin Faller for the NaCl dataset.
+Competing Interests The author declares no Competing Non-Financial Interests but the following Competing Financial Interests: BC has an equity stake in AIMATX Inc.
+Author Contributions BC designed and performed the study, and wrote the paper.